# CosmoBench: A Multiscale, Multiview, Multitask Cosmology Benchmark for Geometric Deep Learning

### v1

@@ 1 Introduction
Cosmological simulations are powerful tools to model the distribution of dark matter and galaxies in the Universe. Astrophysicists use them as virtual laboratories to test hypotheses about the laws and constituents of our Universe. These simulations provide a wealth of data in the forms of point clouds and merger trees. Such data is ripe to be analyzed by methods in geometric deep learning [1], but cosmologists have yet to observe the same breakthroughs that these methods have produced in areas such as computer vision [2], structural biology [3, 4], and climate science [5]. In these areas, rapid progress has been driven by large-scale benchmarks that provide a unified interface to study multiple related tasks using machine learning (ML). Cosmology stands to benefit similarly, as unified benchmarks can facilitate the exchange of ideas between cosmology and ML, driving the development of novel methods and deeper insights into the Universe.
For these reasons, we introduce CosmoBench, a benchmark for geometric deep learning curated from state-of-the-art cosmological simulations. In total, these simulations required more than 41 million core-hours and generated over two petabytes of data. CosmoBench is currently the largest multiscale and multiview benchmark in cosmology: it contains 34 thousand point clouds across three different spatial scales, as well as 25 thousand directed trees on two different temporal scales. Table 1 provides an overview of all the datasets in CosmoBench, and Fig. 1 provides an illustration of the data generation process. We briefly describe each of these sources of data (see Sec. A.1 for a glossary of cosmology terms).
@@
The goal of CosmoBench is to unlock the potential of ML in cosmology. With this goal in mind, CosmoBench consolidates multiple tasks of cosmological interest for ML at scale. These tasks include the prediction of cosmological parameters from point clouds and merger trees (graph-level regression), the prediction of the velocities of individual halos or galaxies from their collective positions (node regression), and the reconstruction of fine-scale merger trees from those on coarser time scales (graph super-resolution). If ML can solve these tasks, then cosmologists will be able to extract more information from observational data, which will improve our knowledge on the fundamental laws and constituents of our Universe.
@@ 2 Related Work
Geometric Deep learning for Cosmology Geometric deep learning has emerged as a promising tool in cosmology, since the cosmic web, along with the dark matter halos and galaxies embedded within it, can be more efficiently described as point clouds than regular grids. This has motivated extensive work in learning accurate mappings from point clouds representing galaxies to cosmological parameters [21, 22, 23, 24, 25, 26, 27, 28], accelerating and improving the necessary modeling of galaxies [29, 30, 31, 32, 33, 34], optimizing observations of the large scale structure of the Universe [35, 36], or reconstructing strong lensing signals [37]. However, recent works also show that off-the-shelf methods from geometric deep learning sometimes struggle with various tasks, such as predicting cosmological parameters solely from point cloud positions [23, 38]. To systematically evaluate deep learning methods, we introduce baselines from cosmology and simple linear models built on interpretable, symmetry-constrained features.
Benchmarks on Point Clouds Popular benchmarks on point clouds mostly come from computer vision (e.g., ShapeNet [39], ModelNet [40]) and structural biology (e.g., AlphaFold DB [3], MoleculeNet [41], Atom3D [42]). Despite the abundance of point clouds from cosmological simulations [6, 7, 8], there is no unified benchmark that formalizes cosmological tasks for ML at scale. A notable recent effort by Balla et al. [43] introduced a subset of 3,560 point clouds from Quijote [6] to evaluate equivariant ML models on predicting cosmological parameters and velocities of individual halos. Yet, this benchmark is limited in size, physical scale, data modality, and task variety. Our benchmark significantly expands over Balla et al. [43]: scaling up to 34,000 point clouds across three simulation suites that model the Universe from linear to deeply non-linear scales, extending to halo merger tree data, and providing more diverse tasks and baselines.
Benchmarks on Graphs Point clouds and directed trees—the central objects in CosmoBench—are structured variants of graph data. While there are many large-scale graph benchmarks [44, 45, 46, 47], most of them are limited to applications in biology and social science. Some graph benchmarks also suffer from improper graph construction or unreliable benchmarking protocol, as pointed out by [48]. CosmoBench offers a high-quality graph benchmark for challenging problems in cosmology, with diverse graph structures, tasks, and a wide spectrum of baselines. It responds to the call in [48] for improving graph benchmarks to spur further progress in graph ML.
@@ 3.1 Multiscale Point Cloud Dataset
Quijote. The Quijote Big Sobol Sequence (BSQ) suite comprises 32,768 dark matter-only $N$-body simulations, each with a box size of $1000~{}\mathrm{cMpc}/h$ and dark matter particle mass of $\sim 10^{12}~{}h^{-1}M_{\odot}$, depending on the value of $\Omega_{\rm m}$. The individual simulations are relatively inexpensive to run due to their low mass resolution, though the full suite required over 35 million core-hours. We exclude 16 simulations due to failures in the halo-finding algorithm, leaving 32,752 simulations in the final dataset. In total five cosmological parameters are sampled in Quijote. We will only focus on $\Omega_{\rm m}$ and $\sigma_{8}$: the matter density and amplitude of matter density fluctuations respectively.
@@ 3.2 Merger Tree Dataset
Data Creation and Split We select all merger trees from CAMELS-SAM with the root node (i.e. the present-time halo at redshift $z=0$) having mass larger than $10^{13}~{}M_{\odot}/h$ ($\sim 10^{5}$ dark matter particles) to ensure that the evolution of the root node’s progenitors is sufficiently well resolved. This yields a collection of more than 460,000 trees across 1,000 cosmological parameter combinations. Since some parameter choices produce significantly more large trees or heavier roots, we randomly select 25 trees per simulation with cosmological parameters $\mathbf{y}=(\Omega_{\rm m},\sigma_{8})$. The original merger trees have up to hundred thousands of nodes and consist of long paths. To reduce storage requirement and condense information, we trim the trees by first removing subtrees with all nodes having mass less than $10^{10}~{}M_{\odot}/h$. We then prune each path where nodes with mass less than $3\times 10^{10}~{}M_{\odot}/h$ are removed. See Fig. A.1 for an illustration. This procedure reduces the tree size by a factor of 5-10, while maintaining most of the merger nodes relevant for probing the merger history, and replacing the smooth halo evolution across a path with a single edge. We then remove any trees with less than 100 nodes. The final CS-Trees dataset contains 24,996 trimmed trees, split per their cosmological parameter choices $(600/204/196)$ (same as the CAMELS-SAM point cloud dataset split).
@@
Task and Evaluation For merger trees, we consider a graph-level regression task to predict the cosmological parameters from tree inputs. We also consider a node-level classification task to determine missing merger nodes, motivated by merger tree temporal “super resolution”  [50, 51]. For this task, we select the 200 largest trees from CS-Trees to avoid degeneracy. Specifically, given a (binary) merger tree $\mathcal{T}_{0}$, we coarsen it by masking out all even time steps, effectively reducing the temporal resolution by half. We arrive at a new (connected) tree by connecting the kept ancestor nodes to their next kept nodes, and save the unique post-merger node for each masked merger node in a dictionary $D$. We then binarize the coarsen tree on nodes with in-degree larger than 2, by keeping the top two pre-merger nodes ranked by their masses, resulting in $\mathcal{T}$. Finally, we append virtual nodes to each merger nodes in tree $\mathcal{T}$, and obtain $\mathcal{T}_{c}=(V+\tilde{V},E+\tilde{E},X\cup\mathbf{0})$. The virtual nodes represent the potentially unresolved mergers. We label a virtual node $\tilde{v}\in\tilde{V}$ as positive if its post-merger node is in the dictionary $D$, and negative otherwise. See Sec. A.3 for details of the full procedure. The node classification task builds a binary classifier for the all virtual nodes given the coarsened (binary) tree. We split the 200 trees randomly into 120/40/40 training/validation/test sets.
@@ 4.1 Predicting Cosmological Parameters from Point Clouds
Graph Neural Networks To capture higher-order clustering information in point clouds, graph neural networks (GNNs) have recently shown encouraging performance for cosmology prediction [62, 63]. A point cloud naturally induces an Euclidean graph by linking pairs of points within some cutoff radius $R_{c}$, also known as a radius graph. We use the point positions as the node features $\mathbf{X}\in{\mathbb{R}}^{n\times 3}$. To respect the underlying translation and reflection symmetries of the cosmological point cloud, we decorate each edge with $3$ features: the normalized pairwise distance $d_{ij}/R_{c}$, and the two dot products $\langle\mathbf{x}_{i},\mathbf{x}_{j}\rangle,\langle\mathbf{x}_{i},(\mathbf{x}_{i}-\mathbf{x}_{j})\rangle$. To go beyond pairwise operation, we identify edge neighbors within the same tetrahedron from a 3-dimensional Delaunay triangulation. We also extract E(3)-invariant features from neighboring node-node, node-edge, and edge-edge pairs using Euclidean and Hausdorff distances, denoted as $\text{Inv}(\cdot,\cdot)$
@@
Here, $\mathcal{N}_{0}(x)$ denotes the neighboring nodes, and $\mathcal{N}_{1}(x)$ the neighboring edges. We use $\bigoplus$ for intra-neighborhood aggregation function and $\otimes$ the inter-neighborhood aggregation function; both are permutation-invariant. Finally, $\psi_{y,z}$ denotes a non-linear updated function determined by the types of $y,z$, and $\sigma$ is an activation function. Our message-passing design is inspired by [64], which supports extensibility to neural networks operating on combinatorial complex topologies—a direction beyond the scope of our benchmarking and reserved for future work. We also perform an ablation on GNNs by removing edge-edge message-passing (i.e., w/o edgeMP). Full details on the network architecture and experiments are deferred to Sec. B.1.
Results and Analysis Table 2 reports the baseline results for predicting $\Omega_{\rm m}$ and $\sigma_{8}$. All of the above methods perform well on the large-scale point clouds in Quijote. In the higher-resolution but smaller-volume CAMELS-SAM simulations, the methods perform worse at predicting $\Omega_{\rm m}$ but similarly or better at predicting $\sigma_{8}$. In CAMELS, which probes the most non-linear scales, the predictions of $\Omega_{\rm m}$ are comparable to those in Quijote, but predictions of $\sigma_{8}$ are severely degraded. This is consistent with the expectation that $\sigma_{8}$ primarily affects rare, high-mass halos, of which only a few are present in CAMELS due to its small volume. Notably, the lightweight LLS model using pairwise statistics achieves performance competitive with GNNs requiring orders of magnitude more parameters and compute time. We also find no significant difference when disabling edge-edge message-passing in GNNs. We remark that a significant body of work has been carried out using the Quijote simulations [^1] https://quijote-simulations.readthedocs.io/en/latest/publications.html [6] to quantify the sensitivity of different summary statistics to $\sigma_{8}$. However, many of these were derived using Fisher matrix calculations and therefore we cannot perform an apple-to-apple comparison.
@@
$\mathbf{R^{2}}$uparrow Quijote CAMELS-SAM CAMELS
@@ 4.2 Predicting Velocities from Positions
where the constant of proportionality is dependent on the cosmological parameters. To provide a cosmology “oracle” for the benchmark, we solve this equation assuming perfect knowledge of cosmological parameters by fitting for the proportionality constant. We solve for the velocity field in Fourier space; further details are provided in Sec. B.2. This solution quantifies the most optimistic possible result achievable using linear theory, whose descriptions begin to fail at small scales. This limitation motivates moving beyond linear theory, for example using nonlinear theory with Bayesian inference [66, 67]; here we explore methods rooted in ML.
@@
where $\phi,\psi$ are MLPs. Further details are reported in Sec. B.2.
Results and Analysis The baseline results for velocity prediction are reported in Tab. 3. All baselines perform relatively well at large scales (Quijote) and deteriorate at smaller scales (CAMELS-SAM and CAMELS). Due to the large number of points in each point cloud, we omit the bootstrap uncertainties on $R^{2}$, as they are negligible. In Quijote, a simple LLS model with only a few parameters outperforms a significantly larger and more computationally expensive GNN. Both LLS and GNN models surpass the linear theory oracle, likely by capturing non-linear corrections to the velocities. Notably, while linear theory requires knowledge of the cosmological parameters, ML methods do not and still outperform it. The GNN is more effective than LLS in CAMELS-SAM, while in CAMELS, linear theory oracle surpasses both. In Fig. 2 we show the true and predicted projected velocity field in a single Quijote point cloud. We provide more visualizations and discussions in Sec. B.2.
@@
$\mathbf{R^{2}}$uparrow Quijote CAMELS-SAM CAMELS

### v2

@@ 1 Introduction
Cosmological simulations are powerful tools to model the distribution of dark matter and galaxies in the Universe. Astrophysicists use them as virtual laboratories to test hypotheses about the laws and constituents of our Universe—such as cosmic expansion and galaxy formation—which cannot be answered through observations alone. These simulations provide a wealth of data in the forms of point clouds and merger trees. Such data is ripe to be analyzed by methods in geometric deep learning [1], but cosmologists have yet to observe the same breakthroughs that these methods have produced in areas such as computer vision [2], structural biology [3, 4], and climate science [5]. In these areas, rapid progress has been driven by large-scale benchmarks that provide a unified interface to study multiple related tasks using machine learning (ML). Cosmology stands to benefit similarly, as unified benchmarks can facilitate the exchange of ideas between cosmology and ML, driving the development of novel methods and deeper insights into the Universe.
For these reasons, we introduce CosmoBench, a benchmark for geometric deep learning curated from state-of-the-art cosmological simulations. In total, these simulations required more than 41 million core-hours and generated over two petabytes of data. CosmoBench is currently the largest multiscale and multiview benchmark in cosmology: it contains 34 thousand point clouds across three different spatial scales, as well as 25 thousand directed trees on two different temporal scales. Table˜1 provides an overview of all the datasets in CosmoBench, and Fig.˜1 provides an illustration of the data generation process. We briefly describe each of these sources of data (see Sec.˜A.1 for a glossary of cosmology terms).
@@
The goal of CosmoBench is to unlock the potential of ML in cosmology. With this goal in mind, CosmoBench consolidates multiple tasks of cosmological interest for ML at scale. These tasks include the prediction of cosmological parameters from point clouds and merger trees (graph-level regression), the prediction of the velocities of individual halos or galaxies from their collective positions (node regression), and the reconstruction of fine-scale merger trees from those on coarser time scales (graph super-resolution). If ML can solve these tasks, then cosmologists will be able to extract more information from simulations and observational data, and from this information they will learn more about the fundamental laws and constituents of our Universe.
@@ 2 Related Work
Geometric Deep learning for Cosmology Geometric deep learning has emerged as a promising tool in cosmology, since the cosmic web, along with the dark matter halos and galaxies embedded within it, can be more efficiently described as point clouds than regular grids. This has motivated extensive work in learning accurate mappings from point clouds representing galaxy distributions to cosmological parameters [21, 22, 23, 24, 25, 26, 27, 28], accelerating and improving the necessary modeling of galaxies [29, 30, 31, 32, 33, 34], optimizing observations of the large scale structure of the Universe [35, 36], or reconstructing strong lensing signals [37]. However, recent works also show that off-the-shelf methods from geometric deep learning sometimes struggle with various tasks, such as predicting cosmological parameters solely from point cloud positions [23, 38]. To systematically evaluate deep learning methods, we introduce baselines from cosmology and simple linear models built on interpretable, symmetry-constrained features.
Benchmarks on Point Clouds Popular benchmarks on point clouds mostly come from computer vision (e.g., ShapeNet [39], ModelNet [40]) and structural biology (e.g., AlphaFold DB [3], MoleculeNet [41], Atom3D [42]). Despite the abundance of point clouds from cosmological simulations [6, 7, 8], there is no unified benchmark that formalizes cosmological tasks for ML at scale. A notable recent effort by Balla et al. [43] introduced a subset of 3,560 point clouds from Quijote [6] to evaluate equivariant ML models on predicting cosmological parameters and velocities of individual halos. Yet this benchmark is limited in size, physical scale, data modality, and task variety. Our benchmark significantly expands over Balla et al. [43] by providing (i) more than 34 thousand point clouds across three simulation suites that model the Universe from linear to deeply non-linear scales, (ii) halo merger tree data, and (iii) more diverse tasks and baselines.
Benchmarks on Graphs Point clouds and directed trees—the central objects in CosmoBench—are structured variants of graph data. While there are many large-scale graph benchmarks [44, 45, 46, 47], most of them are limited to applications in biology and social science. Some graph benchmarks also suffer from improper graph construction or unreliable benchmarking protocol, as pointed out by [48]. CosmoBench offers a high-quality graph benchmark for challenging problems in cosmology, with varying graph structures, diverse tasks, and a wide spectrum of baselines. It responds to the call in [48] for improving graph benchmarks to spur further progress in graph ML.
@@ 3.1 Multiscale Point Cloud Dataset
Quijote. The Quijote Big Sobol Sequence (BSQ) suite comprises 32,768 dark matter-only $N$-body simulations, each with a box size of $1000~\mathrm{cMpc}/h$ and dark matter particle mass of $\sim 10^{12}~h^{-1}M_{\odot}$, depending on the value of $\Omega_{\rm m}$. The individual simulations are relatively inexpensive to run due to their low mass resolution, though the full suite required over 35 million core-hours. We exclude 16 simulations due to failures in the halo-finding algorithm, leaving 32,752 simulations in the final dataset. In total five cosmological parameters ($\Omega_{\rm m},\sigma_{8},\Omega_{\rm b},h,n_{s}$) are sampled in Quijote.
@@ 3.2 Merger Tree Dataset
Data Creation and Split We select all merger trees from CAMELS-SAM with the root node (i.e. the present-time halo at redshift $z=0$) having mass larger than $10^{13}~M_{\odot}/h$ ($\sim 10^{5}$ dark matter particles) to ensure that the evolution of the root node’s progenitors is sufficiently well resolved. This yields a collection of more than 460,000 trees across 1,000 cosmological parameter combinations. Since some parameter choices produce significantly more large trees or heavier roots, we randomly select 25 trees per simulation with cosmological parameters $\mathbf{y}=(\Omega_{\rm m},\sigma_{8})$. The original merger trees have up to hundred thousands of nodes and consist of long paths. To reduce storage requirement and condense information, we trim the trees by first removing subtrees with all nodes having mass less than $10^{10}~M_{\odot}/h$. We then prune each path where nodes with mass less than $3\times 10^{10}~M_{\odot}/h$ are removed. See Fig.˜A.1 for an illustration. This procedure reduces the tree size by a factor of 5-10 and prevent information leakage [^1] The simulation starts with particles whose mass is a simple function of the cosmological parameters; a halo is tracked once $20$ particles are clustered close enough. Thus the ML model can “cheat” by focusing on halos of the smallest mass to predict cosmological parameters, instead of leveraging the halo point cloud structure., while maintaining most of the merger nodes relevant for probing the merger history, replacing the smooth halo evolution across a path with a single edge. We then remove any trees with less than 100 nodes. The final CS-Trees dataset contains 24,996 trimmed trees, split per their cosmological parameter choices $(600/204/196)$ (same as the CAMELS-SAM point cloud dataset split).
@@
Task and Evaluation For merger trees, we consider a graph-level regression task to predict the cosmological parameters from tree inputs. We also consider a node-level classification task to determine missing merger nodes, motivated by merger tree temporal “super resolution”  [50, 51]. For this task, we select the 200 largest trees from CS-Trees to avoid degeneracy. Specifically, given a (binary) merger tree $\mathcal{T}_{0}$, we coarsen it by masking out all even time steps, effectively reducing the temporal resolution by half. We arrive at a new (connected) tree by connecting the kept ancestor nodes to their next kept nodes, and save the unique post-merger node for each masked merger node in a dictionary $D$. We then binarize the coarsen tree on nodes with in-degree larger than 2, by keeping the top two pre-merger nodes ranked by their masses, resulting in $\mathcal{T}$. Finally, we append virtual nodes to each merger nodes in tree $\mathcal{T}$, and obtain $\mathcal{T}_{c}=(V+\tilde{V},E+\tilde{E},X\cup\mathbf{0})$. The virtual nodes represent the potentially unresolved mergers. We label a virtual node $\tilde{v}\in\tilde{V}$ as positive if its post-merger node is in the dictionary $D$, and negative otherwise. See Sec.˜A.3 for details of the full procedure. The node classification task builds a binary classifier for the all virtual nodes given the coarsened (binary) tree. We split the 200 trees randomly into 120/40/40 training/validation/test sets.
@@ 4.1 Predicting Cosmological Parameters from Point Clouds
Graph Neural Networks To capture higher-order clustering information in point clouds, graph neural networks (GNNs) have recently shown encouraging performance for cosmology prediction [62, 63]. A point cloud naturally induces an Euclidean graph by linking pairs of points within some cutoff radius $R_{c}$, also known as a radius graph. We use the point positions as the node features $\mathbf{X}\in{\mathbb{R}}^{n\times 3}$. To respect the underlying translation and reflection symmetries of the cosmological point cloud, we decorate each edge with $3$ features: the normalized pairwise distance $d_{ij}/R_{c}$, and the two dot products $\langle\mathbf{x}_{i},\mathbf{x}_{j}\rangle,\langle\mathbf{x}_{i},(\mathbf{x}_{i}-\mathbf{x}_{j})\rangle$. To go beyond pairwise operation, we identify edge neighbors within the same tetrahedron from a 3-dimensional Delaunay triangulation. We also extract E(3)-invariant features from neighboring node-node, node-edge, and edge-edge pairs using Euclidean and Hausdorff distances, denoted as $\text{Inv}(\cdot,\cdot)$ .
@@
Here, $\mathcal{N}_{0}(x)$ denotes the neighboring nodes, and $\mathcal{N}_{1}(x)$ the neighboring edges. We use $\bigoplus$ for intra-neighborhood aggregation function and $\otimes$ the inter-neighborhood aggregation function; both are permutation-invariant. Finally, $\psi_{y,z}$ denotes a non-linear update function determined by the types of $y,z$, and $\sigma$ is an activation function. Our message-passing design is inspired by [64], which supports extensibility to neural networks operating on combinatorial complex topologies—a direction beyond the scope of our benchmarking and reserved for future work. We also perform an ablation on GNNs by removing edge-edge message-passing (denoted as “w/o edgeMP”). Full details on the network architecture and experiments are deferred to Sec.˜B.1.
Results and Analysis Table˜2 reports the baseline results for predicting $\Omega_{\rm m}$ and $\sigma_{8}$. All of the above methods perform well on the large-scale point clouds in Quijote. In the higher-resolution but smaller-volume CAMELS-SAM simulations, the methods perform worse at predicting $\Omega_{\rm m}$ but similarly or better at predicting $\sigma_{8}$. In CAMELS, which probes the most non-linear scales, the predictions of $\Omega_{\rm m}$ are comparable to those in Quijote, but predictions of $\sigma_{8}$ are severely degraded. This is consistent with the expectation that $\sigma_{8}$ primarily affects rare, high-mass halos, of which only a few are present in CAMELS due to its small volume. Notably, the lightweight LLS model using pairwise statistics achieves performance competitive with GNNs requiring orders of magnitude more parameters and compute time. We also find no significant difference when disabling edge-edge message-passing in GNNs. We remark that a significant body of work has been carried out using the Quijote simulations [^2] https://quijote-simulations.readthedocs.io/en/latest/publications.html [6] to quantify the sensitivity of different summary statistics to $\sigma_{8}$. However, many of these were derived using Fisher matrix calculations and therefore we cannot perform an apple-to-apple comparison. In Sec.˜B.1, we include further ablation studies of 2PCF, LLS, and GNN, as well as the performance of these baselines for predicting three other cosmological parameters in Quijote to illustrate broader applicability yet increasing difficulty in constraining these parameters.
@@
$\mathbf{R^{2}}$ $\uparrow$ Quijote CAMELS-SAM CAMELS
@@ 4.2 Predicting Velocities from Positions
where the constant of proportionality is dependent on the cosmological parameters. To provide a cosmology “oracle” for the benchmark, we solve this equation assuming perfect knowledge of cosmological parameters by fitting for the proportionality constant. We solve for the velocity field in Fourier space; further details are provided in Sec.˜B.2. This solution quantifies the most optimistic possible result achievable using linear theory, whose descriptions begin to fail at small scales. This limitation motivates moving beyond linear theory, for example using nonlinear theory with Bayesian inference [66, 67]; here we explore methods rooted in ML.
@@
where $\phi,\psi$ are MLPs. Further details are reported in Sec.˜B.2.
Results and Analysis The baseline results for velocity prediction are reported in Tab.˜3. All baselines perform relatively well at large scales (Quijote) and deteriorate at smaller scales (CAMELS-SAM and CAMELS). Due to the large number of points in each point cloud, we omit the bootstrap uncertainties on $R^{2}$, as they are negligible. In Quijote, a simple LLS model with only a few parameters outperforms a significantly larger and more computationally expensive GNN. Both LLS and GNN models surpass the linear theory oracle, likely by capturing non-linear corrections to the velocities. Notably, while linear theory requires knowledge of the cosmological parameters, ML methods do not and still outperform it. The GNN is more effective than LLS in CAMELS-SAM, while in CAMELS, linear theory oracle surpasses both. In Fig.˜2 we show the true and predicted projected velocity field in a single Quijote point cloud. We provide more visualizations, ablation studies, and discussions in Sec.˜B.2.
@@
$\mathbf{R^{2}}$ $\uparrow$ Quijote CAMELS-SAM CAMELS
@@
4.2.1 Predicting Velocities from Redshift Positions
Earlier in this section, we assumed access to the real-space positions, representing an idealized case. In practice, however, galaxy positions are measured in so-called redshift space, where the observed galaxy distances are shifted by an amount that depends on their peculiar velocities along the observer’s line of sight. To bridge the gap between simulation data and real observations, we introduce a variant of our velocity prediction task that directly takes inputs as redshift positions. For simplicity, we align the line of sight with the $z$-axis and replace the real-space positions $\{(x_{i},y_{i},z_{i})\}_{i=1}^{n}$ with their redshift counterparts $\{(x_{i},y_{i},\tilde{z}_{i})\}_{i=1}^{n}$, where $\tilde{z}_{i}=z_{i}\cdot(1+\frac{v_{i,z}}{H_{0}})$, $v_{i,z}$ denotes the $z$-component of the velocity and $H_{0}=100~h\,\mathrm{km}\,\mathrm{s}^{-1}\,\mathrm{Mpc}^{-1}$ denotes the “standardized” Hubble constant (since the galaxy positions are expressed in $h^{-1}\,\mathrm{Mpc}$). Note that the redshift effect is more notable at smaller scales. See more details and visualizations in Sec.˜B.3.
As shown in Tab.˜4, replacing real-space positions with redshift-space positions leads to moderately lower performance for the original linear theory and LLS in Quijote, and substantial degradation at smaller scales. To mitigate such degradation, we propose simple modifications of these baselines (denoted as “Modified”), essentially shrinking the line-of-sight distance to counteract the redshift effect, which shows notable improvement in CAMELS-SAM and CAMELS; see details in Sec.˜B.3. In contrast, while the GNN exhibits similar performance drop to other baselines in Quijote, it performs better in CAMELS-SAM and CAMELS using redshift position inputs. This arises from GNN’s ability to leverage the line-of-sight velocity information encoded in the redshift positions, effectively turning the original regression problem into a simpler denoising task, as evidenced by the per-axis performance reported in Tab.˜11.
Table 4: Velocity Prediction from Redshift Positions. Numbers marked with a denote that additional “cosmological oracle” information was used. For linear theory and LLS, each dataset column is split into results from the original baseline (see Sec.˜4.2) and modified variants (see Sec.˜B.3).
Quijote CAMELS-SAM CAMELS
$\mathbf{R^{2}}$ $\uparrow$ Original Modified Original Modified Original Modified
Linear theory oracle 0.3269 0.3269 0.1177 0.1621 0.0615 0.0732
LLS 0.3367 0.3362 0.1172 0.2116 0.0801 0.2185
GNN 0.3500 0.3177 0.3197